# Deep Learning for Pathological Speech: A Survey

### v1

@@ Abstract
Advancements in spoken language technologies for neurodegenerative speech disorders are crucial for meeting both clinical and technological needs. This overview paper is vital for advancing the field, as it presents a comprehensive review of state-of-the-art methods in pathological speech detection, automatic speech recognition, pathological speech intelligibility enhancement, intelligibility and severity assessment, and data augmentation approaches for pathological speech. It also highlights key challenges, such as ensuring robustness, privacy, and interpretability. The paper concludes by exploring promising future directions, including the adoption of multimodal approaches and the integration of graph neural networks and large language models to further advance speech technology for neurodegenerative speech disorders.
@@ I Introduction
Speech production is a complex mechanism involving cognitive planning, coordinated muscle activity, and sound creation [1]. The process starts in the brain with the conceptualization of a message, followed by the organization of phonetic and prosodic plans such as rhythm and style. The motor cortex then orchestrates the activation of approximately $100100$ muscles, enabling the articulatory organs such as the tongue, lips, and jaw, to shape the vocal tract and produce specific sounds. The initial sound is generated in the larynx, where air from the lungs causes the vocal folds to vibrate. These phonatory structures adjust voice quality and prosody, while the articulatory organs further refine the sound by altering the shape of the vocal tract. Finally, the resulting speech is emitted through the oral and nasal cavities. Given the intricate coordination required for this process, any disruption to these finely tuned mechanisms can severely impair communication.
@@
Neurodegenerative disorders such as Parksinson’s disease (PD), Amyotrophic Lateral Sclerosis (ALS), or Alzheimer’s disease are leading causes of such disruptions [2, 3, 4]. These disorders impair the brain regions and motor systems responsible for initiating, planning, and controlling the movements needed for speech production, resulting in a variety of speech disorders such as dysarthria, aphasia, apraxia of speech, dysphonia, or stuttering [5, 6, 7, 4, 8, 9]. Dysarthria and apraxia of speech, commonly seen in PD and ALS, are characterized by articulation deficiencies, vowel distortions, reduced loudness variation, hypernasality, or syllabification [5, 4]. Dysphonia, also frequent in PD and ALS, is marked by abnormal voice quality such as hoarseness and breathiness [6]. In contrast, aphasia typically presents as difficulties with word-finding, sentence construction, and understanding spoken or written language, and is most commonly associated with Alzheimer’s disease or other forms of dementia [8]. Lastly, stuttering, characterized by disruptions in the flow of speech such as repetitions or prolongations can also arise in neurodegenerative disorders like Parkinson’s disease [9].
@@
Because of increasing population numbers and aging, the prevalence of neurological disorders, and hence, of these various speech disorders, is also rapidly increasing. In $20192019$, the World Health Organization estimated that over $8.58.5$ million people worldwide were living with PD [10], up from $6.16.1$ million in $20162016$ [11] and $2.52.5$ million in $19901990$ [11]. By $20402040$, this figure is projected to surpass $1717$ million [11]. Similarly, dementia affected more than $4646$ million people globally in $20152015$ and this figure is expected to rise to $131.5131.5$ million by $20502050$ [12]. The prevalence of ALS is also growing significantly, with cases anticipated to increase by nearly 70% between $20152015$ and $20402040$ [13]. This increasing prevalence of neurological disorders, and consequently of the associated speech disorders, underscores the need to prioritize speech disorders both in the context of clinical practice as well as in the context of speech-based technologies.
In clinical practice, diagnosing the presence of a speech disorder (i.e., discriminating between neurotypical and impaired speech) is crucial, since the presence of a disorder can be one of the earliest signs of neurodegenerative disorders [14, 15, 16]. Further, an accurate differential diagnosis of the speech disorder (e.g., discriminating between dysarthria and apraxia of speech) is also important, since it can provide clues about the underlying neuropathology [17, 18]. Additionally, monitoring speech characteristics such as severitys and intelligibility after diagnosis is essential for tracking disease progression and assessing the effectiveness of speech therapy interventions over time [2, 19]. Traditionally, clinicians conduct such evaluations through costly and time-consuming auditory-perceptual assessments as depicted in Fig. 2. However, because of the difficulty of detecting clinical-perceptual characteristics by ear (particularly in the presence of mild impairments) and because the clinical-perceptual characteristics of several speech disorders (e.g., of dysarthria and apraxia of speech) overlap, diagnosing the presence of a disorder and discriminating between various disorders is hard for non-expert clinicians and even expert inter-rater agreement can be low [20, 21]. In the context of speech-based technologies, patients with speech disorders often struggle to access and effectively use various virtual assistant tools like Cortana, Alexa, and Siri [22]. As speech disorders become more prevalent with neurodegenerative conditions, it is critical to prioritize both the diagnosis and treatment of these disorders in clinical practice, while also ensuring that patients with such impairments have equitable access to speech-based technologies. Addressing these barriers could lessen the burden on the health care system and significantly improve the patients’ quality of life and their ability to engage with everyday digital tools.
Figure 1: Traditional auditory-perceptual assessment in clinical practice. The clinician listens to the (potential) patient speaker and assesses by ear the various characteristics of the speech.
@@
Figure 2: Automatic pathological speech analysis system. The automatic model is trained to detect and analyze speech impairments. Clinicians may use the insights provided by the automatic model to organize therapeutic sessions accordingly. Additionally, they may perform manual acoustic analysis to gain further insights into the patient’s speech patterns and impairments, providing complementary information to the automatic model’s decision.
@@
Aiming to assist the clinical diagnosis and treatment of patients suffering from neurodegenerative disorders, there has been a growing interest in the research community to develop automated and objective methods for pathological speech analysis as shown in Figure 2. These advanced technologies are designed to minimize bias, enhance diagnostic accuracy, and streamline the assessment process, ensuring greater efficiency and consistency. Clinicians can then use insights provided by the automatic models to organize therapeutic sessions accordingly, ensuring that the treatment addresses the specific impairments identified. Further, clinicians may perform additional manual acoustic analysis to gain further insights into the patient’s speech patterns and impairments, providing complementary information to the automatic model’s decision for a comprehensive therapeutic approach. Besides the clinical domain, efforts have been directed towards developing various speech-based technological applications aimed at pathological speakers, such as automatic speech recognition systems (ASRs) [23], speech synthesis systems [24, 25, 26, 27] or intelligibility enhancement solutions [28, 29, 30, 31]. To the best of our knowledge, there is currently no comprehensive survey paper discussing the research directions and challenges of this area both from a clinical as well as a technological perspective. Although there have been related works such as [32], they primarily provide an overview of acoustic-articulatory features in neurodegenerative disorders. Other studies such as [33, 34, 35, 36, 37] mainly focus on the discrimination between neurotypical and impaired speech. In $20162016$, [38] discussed some of the wider challenges faced in the pathological speech domain. However, this work addresses only a limited set of challenges and is now somewhat outdated, given the rapid advancements in the field over the past decade. In contrast, as depicted in Figure 3, our work aims to fill this gap by providing an extensive review of the field encompassing pathological speech from various clinical and technological perspectives, such as automatic discrimination between neurotypical speech and speech disorders, ASR systems for pathological target speakers, enhancement systems aiming to enhance the intelligibility of pathological speech, severity and intelligibility assessment and data augmentation approaches.
Figure 3: Overview of Deep Learning for Pathological Speech: This framework leverages various input features and datasets processed through deep learning models for tasks including detection, recognition, enhancement, assessment and explainaibility of pathological speech. It employs a range of metrics to evaluate model performance, providing a comprehensive approach to improving speech impairment detection and treatment.
@@
The remainder of the paper is organized as follows. Section II provides a high level overview of the handcrafted acoustic features used in pathological speech analysis. Section III describes various accessible and non accessible pathological datasets employed in the literature. Section IV describes the different approaches to automatic pathological speech detection. Section V examines pathological speech in the context of ASR systems. Section VI discusses speech enhancement techniques aimed at improving pathological speech intelligibility. Section VII describes approaches for automatically estimating the severity and intelligibility of pathological speakers. Section VIII summarizes the various data augmentation methods used in automatic pathological speech systems. The robustness issues arising in this field are discussed in Section IX-A. Finally, Section X explores the interpretability of pathological speech models. In summary, our contributions are the following:
@@
    We present the first comprehensive survey of automatic approaches for pathological speech from a clinical and technological perspectives ranging from detection, recognition, enhancement, assessment and interpretability.
@@
II Pathological Speech: Acoustic Characteristics and Representation
Pathological speech exhibits a range of acoustic anomalies, including deviations in pitch, loudness, vowel space reduction, and articulation [39, 40]. Additionally, it can lead to asymmetrical tension in the vocal folds, resulting in irregular vibrations and, consequently, an abnormal fundamental frequency. Excessively high or low fundamental frequency, combined with excessive vocal intensity, can exacerbate the severity of pathological voice conditions, producing characteristics such as shrillness, screechiness, hoarseness, or huskiness [41]. Moreover, patients (such as those diagnosed with PD) can also show inconsistent rhythmic structures in comparison to control groups under various auditory feedback conditions [42]
To capture such abnormalities as biomarkers for automatic pathological speech detection, researchers have employed various handcrafted acoustic features such as OpenSMILE [43], Mel-frequency cepstral coefficients (MFCCs) [44], or spectro-temporal sparsity features [45]. Further, various time-frequency representations such as the short-time Fourier transform (STFT) or the Mel spectrogram have also been directly exploited in combination with deep learning approaches to directly learn pathology-discriminant cues [46]. More recently, latent embeddings derived from self-supervised models have been used as more powerful representations of speech patterns and the various impairments [43, 47, 48]. It is important to note that a significantly larger set of features and representations have been employed to characterize disorders across various pathologies than the ones outlined above. However, here we focus on highlighting the most commonly utilized ones, which will be further elaborated upon in the following.
II-A Handcrafted Acoustic Features
OpenSMILE
Among the various features used in pathological speech detection, OpenSMILE features have been widely explored in the literature in conjunction with traditional machine learning algorithms [49, 50, 51, 52, 31, 53, 54, 55, 56, 57, 58, 59, 60, 61]. The OpenSMILE feature set includes a 6552-dimensional feature vector that primarily consists of low-level audio features such as CHROMA, CENS, loudness, MFCCs, and other spectral features. In the context of pathological speech, these features can capture the subtle anomalies in speech patterns that are indicative of disorders. OpenSMILE’s comprehensive feature set allows for the analysis of prosodic elements like pitch, jitter, shimmer, and formant frequencies, which are crucial for identifying and differentiating various speech pathologies. However, OpenSMILE features are very general features that have been used for a variety of speech applications such as emotion recognition [62], speech recognition [63] and they are not specifically handcrafted to capture pathological cues, thus may not be the most suitable features for pathological speech analysis.
MFCCs
MFCCs have garnered significant attention in pathological speech detection due to their potential to characterize articulation deficiencies [64, 65, 66, 67, 68, 69, 70, 71, 72]. While MFCCs have demonstrated promising results, they exhibit high sensitivity to background noise and present challenges in real-time applications due to their complex nature and lack of interpretability.
Spectro-temporal sparsity
II-B Time-Frequency Representations
II-C Self Supervised Embeddings
Even though handcrafted acoustic features and time-frequency input representations have achieved promising results in the analysis of pathological speech, their performance remains limited. The current state-of-the-art input representations for pathological speech primarily consist of latent embeddings derived from self-supervised models such as wav2vec 2.0 (w2v2) [88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 47, 101, 102, 103, 104, 61, 105, 106, 107, 108, 109, 110, 111]. These advanced models have proven to be more effective in capturing complex speech patterns and nuances, leading to improved performance in tasks like detection and recognition [112, 48], However, there is still ongoing research to further refine these embeddings and explore additional models to enhance their robustness and generalization across diverse pathological speech conditions [113, 114, 115].
III Pathological Speech Datasets
@@
Datasets for pathological speech research remain scarce due to several inherent challenges in data collection. One of the primary difficulties is the sensitive nature of the population involved. Recruiting participants suffering from neurological disorders requires careful consideration of ethical and privacy concerns, as well as navigating the potential stigma associated with such disorders. Additionally, the physical and cognitive challenges faced by these individuals can complicate the process of obtaining high-quality speech recordings. Another key challenge is the variability in pathological speech patterns. Speech impairments can manifest in numerous ways and may vary widely across individuals, even within the same diagnostic category. This variability makes it difficult to create standardized protocols for data collection that ensure consistency and relevance across samples. Moreover, speech impairments may fluctuate over time, further complicating the process of capturing representative speech samples. When collecting datasets for pathological speech research, it is essential to prioritize inclusivity, ensuring a diverse representation of different speech disorders, age groups, and demographics. Additionally, data collection protocols should consider the comfort and cooperation of participants. Ethical considerations must also be at the forefront, ensuring informed consent and the protection of participant privacy. Lastly, it is important to design flexible and scalable collection methods that can capture a range of speech characteristics while maintaining consistent quality across diverse individuals and conditions. In the remainder of this section, we briefly review pathological datasets commonly used in the literature. Please note that due to restricted access to most of the pathological neurodegenrative datasets, we are unable to provide detailed information such as duration, number of male and female speakers, etc., for all of them. A summary of these datasets and their characteristics is presented in Table I.
@@
TORGO  [116].  The TORGO dataset contains speech recordings from English speaking subjects (controls and patients) and the corresponding three dimensional articulatory features. The patients suffer from ALS or Cerebral Palsy (CP). The dataset consists of seven (three females, four males) controls and eight (three females, five males) dysarthric speakers respectively.
PC-GITA [117].  The PC-GITA dataset is a Spanish dataset containing speech recordings from $5050$ controls and $5050$ patients suffering from PD. The two groups of speakers are age- and gender-matched, with $2525$ female and $2525$ male speakers in each group.
MoSpeeDi [118].  The MoSpeeDi dataset is a French dataset containing speech recordings from $466466$ controls and $138138$ patients suffering from various types of motor speech disorders such as dysarthria or apraxia of speech. While subgroups of age- and gender-matched controls and patients can be found within the database, the overall dataset is not age- and gender-matched.
Nemours [119].  The Nemours dataset is an English dataset containing speech recordings from $1111$ male speakers with varying degrees of dysarthria severity. This dataset includes a total of 74 sentences and 814 short nonsense sentences.
CUDYS [120].  The Cantonese Dysarthric Speech Corpus (CUDYS) contains speech recordings from $55$ controls and $1111$ dysarthric patients. The control group consists of $22$ female and $33$ male speakers whereas the dysarthric group consists of $55$ female and $66$ male speakers. The dataset has a total duration of $1010$ hours, with the speakers’ ages ranging between $2525$ and $6767$ years.
AMSDC [121].  The Atlanta Motor Speech Disorders Corpus (AMSDC) is an English dataset containing speech recordings from $9999$ ($6262$ males and $3737$ females) patients diagnosed with neurogenic disorders. This dataset is essential for studying motor speech disorders, though access to the full dataset might be restricted.
EST [122].  The EST dataset is a Dutch dataset containing speech recordings from 16 dysarthric patients ($1616$ males). The dataset has a total duration of $\approx$ $66$ hours and the speakers’ ages range between $3434$ and $7575$ years.
EasyCall [123].  The EasyCall dataset is an Italian dataset containing speech recordings from $2424$ controls and $3131$ patients diagnosed with PD, Huntington’s disease, ALS, and peripheral neuropathy. The control group consists of $1010$ female and $1414$ male speakers whereas the patient group consists of $1111$ female and $2020$ male speakers. The corpus includes a total of 21,386 speech commands, with 10,077 recordings from healthy subjects and 11,309 from dysarthric patients.
COPAS [124].  The Corpus of Pathological and Normal Speech (COPAS) contains speech recordings from Dutch speakers with $197197$ pathological speakers and $122122$ normal speakers. The database comprised of 8 distinct pathological categories.
ParkCeleb [125].  The previously reviewed datasets are not longitudinal and do not allow tracking the progression of the speech disorder within the same patient along time. To address this gap, the English ParkCeleb dataset was recently introduced in [125]. This dataset contains speech recordings from $4040$ controls and $4040$ patients suffering from PD.
Summary of datasets.  As outlined in Table I, many of the pathological speech datasets used in the literature are private and not openly accessible, limiting their availability for broader research and replication. Furthermore, datasets such as TORGO, although publicly available, have been shown to contain detrimental recording artefacts [43], which limits their suitability for studying pathological speech characteristics or for developing automatic pathological speech detection approaches [43, 113]. This lack of accessible, high-quality data creates a substantial barrier for developing and validating effective algorithms and tools. Consequently, the advancement in understanding and treating pathological speech conditions is slowed, limiting the potential for innovative solutions and breakthroughs in this critical area of healthcare. Ensuring the availability of clean, reliable, and comprehensive datasets is essential to foster meaningful research and development. Additionally, many existing pathological speech datasets are predominantly in English, restricting their applicability to non-English-speaking populations and limiting cross-linguistic comparisons. This linguistic bias creates challenges when attempting to develop universal speech models or tools that can be applied globally. Furthermore, most of these datasets are relatively small, often comprising only a limited number of participants, which hinders the development of robust, generalizable models. The small size of these datasets also makes it difficult to account for the variability in speech disorders across different populations and conditions. Another important limitation is the imbalance in demographic representation in terms of gender. Many datasets suffer from a skewed representation, often over-representing one gender, which can lead to biases in speech models. Additionally, some datasets may not include a sufficient variety of age groups or pathological conditions, further restricting their generalizability. These factors collectively underscore the need for high quality, larger, more diverse, and well-balanced pathological speech datasets that reflect the complexity and diversity of real-world populations.
Figure 4: Automatic Pathological Speech Detection Pipeline: The process begins with recording the patient’s speech, which is then processed to extract features that are meaningful and suitable for an AI classifier. These features are transformed into a feature space, where a neural network model analyzes them to make a diagnostic decision. This pipeline streamlines the detection process, enabling efficient and accurate identification of pathological speech patterns.
IV Automatic Pathological Speech Detection
TABLE I: Overview of Pathological Speech Databases. (PD: Parkinson’s disease, CVA: Cerebrovascular accident, TBI: Traumatic brain injuries, ALS: Amyotrophic lateral sclerosis, Huntington’s disease, 3D: Articulatory (3D), L: Longitudinal data, Accessible: Accessible through a license for research purposes )
@@ IV Automatic Pathological Speech Detection
TORGO 07 08 ALS/CP English Speech + (3D) Public

### v2

# Overview of Automatic Speech Analysis and Technologies for Neurodegenerative Disorders: Diagnosis and Assistive Applications

@@ Abstract
Advancements in spoken language technologies for neurodegenerative speech disorders are crucial for meeting both clinical and technological needs. This overview paper is vital for advancing the field, as it presents a comprehensive review of state-of-the-art methods in pathological speech detection, automatic speech recognition, pathological speech intelligibility enhancement, intelligibility and severity assessment, and data augmentation approaches for pathological speech. It also highlights key challenges, such as ensuring robustness, privacy, and interpretability. The paper concludes by exploring promising future directions, including the adoption of multimodal approaches and the integration of large language models to further advance speech technologies for neurodegenerative speech disorders.
@@ I Introduction
Speech production is a complex mechanism that involves cognitive planning, coordinated muscle activity, and sound creation [1]. The process starts in the brain with the conceptualization of a message, followed by the organization of phonetic and prosodic plans, such as rhythm and style. The motor cortex then orchestrates the activation of approximately $100$ muscles, allowing articulatory organs such as the tongue, lips, and jaw to shape the vocal tract and produce specific sounds. The initial sound is generated in the larynx, where the air from the lungs causes the vocal folds to vibrate. These phonatory structures adjust voice quality and prosody, while the articulatory organs further refine the sound by altering the shape of the vocal tract. Finally, the resulting speech is emitted through the oral and nasal cavities. Given the intricate coordination required for this process, any disruption of these finely tuned mechanisms can severely alter communication and result in pathological speech.
In this work, we define pathological speech as speech that deviates from neurotypical patterns due to underlying impairments. These deviations can manifest along multiple dimensions, including voice, articulation, prosody, and language. Voice impairments involve abnormalities in vocal fold vibration or in breath control, leading to hoarseness, breathiness, or a strained voice [2, 3]. Articulation impairments involve abnormalities in the coordination or movement of the various articulators, leading to slurred, imprecise, or segmented speech [4, 5]. Furthermore, prosodic impairments involve abnormalities in the rhythm, stress, or intonation of speech, leading to speech that may sound flat or monotonic [6]. Finally, language impairments involve abnormalities in the formulation or comprehension of linguistic content, leading to difficulties with word retrieval, sentence construction, or understanding spoken or written language [7]. While pathological speech can arise from a wide range of neurological, structural, or functional impairments, our objective is to provide an overview of automated methods and speech-based technologies targeting pathological speech arising due to neurodegenerative disorders.
@@
Neurodegenerative disorders such as Parkinson’s disease (PD), Amyotrophic Lateral Sclerosis (ALS), or Alzheimer’s disease are leading causes of voice, articulation, prosody, and language disruptions [8, 9, 4]. These disorders impair the brain regions and motor systems responsible for initiating, planning, and controlling the movements needed for speech production, resulting in a variety of speech disorders such as dysarthria, aphasia, apraxia of speech, or dysphonia [5, 2, 10, 4, 7, 11, 12]. Dysarthria and apraxia of speech, commonly seen in PD and ALS, are primarily characterized by articulatory and prosodic deficiencies [13], vowel distortions, reduced loudness variation, hypernasality, or syllabification [5, 4]. Dysphonia, also frequent in PD and ALS, is marked by abnormal voice quality such as hoarseness and breathiness [2, 3]. In contrast, aphasia typically presents as difficulties with word-finding and sentence construction, and is most commonly associated with Alzheimer’s disease or other forms of dementia [7].
@@
As the population grows and ages, the prevalence of neurological disorders, and consequently of various speech disorders, rapidly increases. In $2019$, the World Health Organization estimated that over $8.5$ million people worldwide were living with PD [14], up from $6.1$ million in $2016$ [15] and $2.5$ million in $1990$ [15]. By $2040$, this figure is projected to surpass $17$ million [15]. Similarly, dementia affected more than $46$ million people globally in $2015$ and this figure is expected to rise to $131.5$ million by $2050$ [16]. The prevalence of ALS is also growing significantly, with cases anticipated to increase by nearly 70% between $2015$ and $2040$ [17]. This increasing prevalence of neurological disorders, and consequently of the associated speech disorders, underscores the need to prioritize speech disorders both in the context of clinical practice as well as in the context of speech-based technologies.
Accurately diagnosing the presence of speech disorders in clinical practice (i.e., distinguishing between neurotypical and impaired speech) is crucial, as the presence of such disorders may serve as an early indicator of neurodegenerative conditions [18, 19, 20]. Further, an accurate differential diagnosis (e.g., discriminating between dysarthria and apraxia of speech) can provide important clues about the underlying neuropathology [21, 22]. Monitoring speech characteristics such as severity and intelligibility after diagnosis is also essential for tracking disease progression and evaluating the effectiveness of speech therapy interventions over time [8, 23, 12, 6].
Speech assessment in clinical practice relies on established perceptual evaluation scales that serve as gold standards for diagnosing and characterizing various aspects of speech impairments. For example, the GRBAS scale [24, 25] evaluates voice quality through parameters like grade, roughness, breathiness, asthenia, and strain. The Consensus Auditory Perceptual Evaluation-Voice (CAPE-V) [23] offers a similar perceptual framework, excluding asthenia. In addition to these general assessment scales, specific scales are employed for specific conditions. For instance, the Unified Parkinson’s Disease Rating Scale (UPDRS) [26] includes components for speech and motor function assessment in PD, while the Bogenhausen Dysarthria Scales (BoDyS) [27] focus on the severity and profile of dysarthric impairments. Intelligibility, a key outcome measure in many disorders, is often assessed using tools such as the Assessment of Intelligibility of Dysarthric Speech [28], which standardizes both single-word and sentence intelligibility evaluations.
Traditionally, clinicians conduct these evaluations through costly and time-consuming auditory-perceptual assessments [12, 23], as illustrated in Fig.1. Diagnosing speech disorders and distinguishing between various conditions can be particularly challenging, even for experienced clinicians. This difficulty arises from the subtle nature of clinical-perceptual characteristics which are often hard to detect by ear, especially in cases of mild impairments. The overlapping characteristics of certain speech disorders, such as dysarthria and apraxia of speech, further complicate the process. Consequently, inter-rater agreement for (differential) diagnosis of speech disorders among clinicians can be low [29, 30].
These clinical challenges highlight the growing need for complementary, technology-driven approaches to support diagnosis, monitoring, and intervention. In parallel, the increasing prevalence of speech disorders poses significant accessibility challenges in patients’ everyday interactions with speech-based technologies. For example, individuals with dysarthria or apraxia of speech often experience difficulty using mainstream virtual assistants such as Cortana, Alexa, and Siri [31]. As speech disorders become more prevalent with neurodegenerative conditions, it is critical to prioritize both the diagnosis and treatment of these disorders in clinical practice, while also ensuring that patients with such impairments have equitable access to speech-based technologies. Addressing these barriers could lessen the burden on the health care system and significantly improve the patients’ quality of life and their ability to engage with everyday digital tools.
@@
Figure 1: Traditional auditory-perceptual assessment in clinical practice (bounded by the dashed box A) and automatic pathological speech analysis system (bounded by the solid box B). The clinician listens to the (potential) patient and assesses by ear the various characteristics of the speech. The automatic model is trained to detect and analyze speech impairments. Clinicians may use the insights provided by the automatic model to organize therapeutic sessions accordingly. Additionally, they may perform manual acoustic analysis to gain further insights into the patient’s speech patterns and impairments, providing complementary information to the automatic model’s decision.
@@
Aiming to assist the clinical diagnosis and treatment of patients suffering from neurodegenerative disorders, there has been a growing interest in the research community to develop automated and objective methods for pathological speech analysis. A schematic illustration of such analysis is depicted in Fig. 1. These advanced technologies are designed to minimize bias, enhance diagnostic accuracy, and streamline the assessment process, ensuring greater efficiency and consistency. Clinicians can then use insights provided by the automatic models to organize therapeutic sessions accordingly, ensuring that the treatment addresses the specific impairments identified. Further, clinicians may perform additional manual acoustic analysis to gain further insights into the patient’s speech patterns and impairments, providing complementary information to the automatic model’s decision for a comprehensive therapeutic approach. Besides the clinical domain, efforts have been directed towards developing various speech-based technological applications aimed at pathological speakers, such as automatic speech recognition systems (ASRs) [32], speech synthesis systems [33, 34, 35, 36] or intelligibility enhancement solutions [37, 38, 39, 40]. To our knowledge, there is currently no comprehensive survey paper that discusses the research directions and challenges of this area both from a clinical and a technological perspective. Although, there have been related works, such as [41, 42, 43], they are primarily focused on specific characteristics, applications, or disorders, limiting their scope. For instance, [44] provides an overview of acoustic-articulatory characteristics in neurodegenerative disorders. Other studies, such as [45, 46, 47, 48, 49, 50], mainly focus their review on the discrimination between neurotypical and impaired speech. Gupta et al. [51] discussed some of the wider challenges faced in the pathological speech domain. However, this work addresses only a limited set of challenges and is now somewhat outdated given the rapid advancements in the field over the past decade. In contrast, as depicted in Fig. 2, our work aims to fill this gap by providing an extensive review of the field encompassing pathological speech from various clinical and technological perspectives, such as automatic discrimination between neurotypical speech and speech disorders, ASR systems for pathological target speakers, enhancement systems aiming to enhance the intelligibility of pathological speech, severity and intelligibility assessment, and data augmentation approaches.
Figure 2: Overview of key components discussed in this manuscript, including datasets, features, and research directions for pathological speech.
@@
The remainder of the paper is organized as follows. Section II describes various pathological datasets employed in the literature. Section III provides a high level overview of the various speech representations used in pathological speech analysis. Section IV describes the different approaches to automatic pathological speech detection. Section V examines pathological speech in the context of ASR systems. Section VI discusses speech enhancement techniques aimed at improving pathological speech intelligibility. Section VII describes approaches for automatically estimating the intelligibility and severity of pathological speakers. Section VIII summarizes the various data augmentation methods used in automatic pathological speech systems. Finally, Section IX presents challenges and promising future research directions in the field. In summary, our contributions are the following:
@@
    We present the first comprehensive survey of automatic approaches for pathological speech from a clinical and technological perspectives ranging from detection, recognition, enhancement, and assessment.
@@
II Pathological Speech Datasets
@@
Datasets for pathological speech research are scarce due to several inherent challenges in data collection. One of the primary difficulties is the sensitive nature of the population involved. Recruiting participants suffering from neurological disorders requires careful consideration of ethical and privacy concerns, as well as navigating the potential stigma associated with such disorders. Additionally, the physical and cognitive challenges faced by these individuals can complicate the process of obtaining high-quality speech recordings. Another key challenge is the variability in pathological speech patterns. Speech impairments can manifest in numerous ways and may vary widely across individuals, even within the same diagnostic category. This variability makes it difficult to create standardized protocols for data collection that ensure consistency and relevance across samples. Moreover, speech impairments may fluctuate over time, further complicating the process of capturing representative speech samples. When collecting datasets for pathological speech research, it is essential to prioritize inclusivity, i.e., ensuring a diverse representation of different speech disorders, age groups, and demographics. Additionally, data collection protocols should consider the comfort and cooperation of participants. Ethical considerations must also be at the forefront, ensuring informed consent and the protection of participant privacy. Lastly, it is important to design flexible and scalable collection methods that can capture a range of speech characteristics while maintaining consistent quality across diverse individuals and conditions. In the remainder of this section, we briefly review pathological datasets [^1] Due to restricted access to many datasets, we are unable to provide complete metadata, including details like the number of male and female speakers commonly used in the literature. The summary of these datasets and their characteristics is presented in Table I.
@@
  • TORGO  [52].  The TORGO dataset contains English (spontaneous and read) speech recordings from control speakers and patients and the corresponding three-dimensional electromagnetic articulography (EMA). The patients suffer from ALS or Cerebral Palsy (CP). The dataset consists of recordings from 77 (33 female, $4$ male) control speakers and 88 (33 female, $5$ male) patients.
    TABLE I: Overview of pathological speech datasets (PD: Parkinson’s disease, CVA: Cerebrovascular accident, TBI: Traumatic brain injuries, ALS: Amyotrophic lateral sclerosis, Huntington’s disease, AD: Alzheimer’s Disease, MCI: Mild Cognitive Impairment, EMA: Electromagnetic Articulography, EGG: Electroglottography, L: Longitudinal data). Public datasets are openly available on the web and can be downloaded without the need for explicit approval or signing an agreement. Accessible datasets require an application process or signing an agreement, but can be obtained by the research community under defined conditions. Non-accessible datasets are private and not available to the wider research community.
@@
    TORGO 07 08 ALS/CP English Speech + EMA Public
@@
    Saarbrucken Voice Database 869 1,356 Pathological German Speech + EGG Public
    Turkish Parkinson Dataset 20 20 PD Turkish Speech Not Accessible
    Czech Parkinson’s Dataset 22 61 PD Czech Speech Public
    Italian Parkinson’s Database 28 37 PD Italian Speech Public
    NeuroVoz 58 54 PD Spanish Speech Public